AttackOnTitanGirl:

In geometric progression, the ratio of each number to the preceding one is the same. Which is an example of a geometric progression? 1, 2, 3 4, 6, 8 3, 21, 35 4, 12, 36

Vocaloid:

for each sequence you need to check if term2/term1 = term3/term2 for example, for option 1, term2/term1 = 2/1 and term3/term2 = 3/2 so this series is not geometric apply this method to every answer choice and select all choices that fulfill this requirement